OpenStudy (anonymous):

Find S8 for the series -2 + -10 + -50 + -250 +…

OpenStudy (anonymous):

sn = a1(1 - r^n) / (1 - r) a1 = -2 r = 5 n = 8 s8 = -2(1 - 5^8) / (1 - 5) s8 = -2( 1 - 390625) / -4 s8 = -2(- 390624) / -4 s8 = 781248/-4 s8 = - 195312